hey my teacher gave us this problem and me and my friend cant figure out how to do it everytime we try we end up getting the wrong answer here it is jeans cost $6 dollars more than shorts. If joan buys 3 pairs of shorts and 2 pairs of jeans for $102, how much does one pair cost?
Answer by checkley71(8403)

J=6.00+S AND 3S+2(6.00+S)=102.00 OR 3S+12.00+2S=102.00 OR 5S+12.00=102.00
OR 5S=102.00-12.00 OR 5S=90.00 OR S=90.00/5 OR SHORTS=$18.00
THEREFORE J=6.00+18.00 OR JEANS=$24.00.
PROOF 3X18.00+2x24.00=102.00 OR 54.00+48.00=102 OR 102.00=102.00.